. For an economy with a tax rate of 10%, the following is given:
C = 60 + 0.8YD
G = 400
I = 140
NX = 10 - 0.20Y
(I) If the economy produces $1 of extra output, how much of it is;
(a). spent on domestic goods?
(b). import expenditure?
(c). saved?
1
Expert's answer
2015-02-23T13:58:05-0500
(I) If the economy produces $1 of extra output, it will:
(a) spent on domestic goods 0.8*$1 = $0.8 more.
(b) spent 0.2*$1 = $0.2 more on imports
(c) as savings S = s*Y, s = 1 - c and c = 0.8 in our case, savings will increase by (1 - 0.8)*$1 = $0.2.